A block has a mass of 12 kg and is accelerating at 2.5 m/s^2 to the right. What is the force being applied to it?
1
Expert's answer
2020-11-16T07:53:11-0500

We can find the force applied to the block from the Newton's Second Law of Motion:


"F=ma=12\ kg\cdot 2.5\ \dfrac{m}{s^2}=30\ N."

Answer:

"F=30\ N."